# Polarization modes of gravitational waves in general symmetric teleparallel gravity

### v1

@@ Abstract
Abstract: This paper investigates the polarization modes of gravitational waves within the most general symmetric teleparallel gravity theory that allows for second-order field equations We consider both scenarios where test particles either carry or do not carry a hypermomentum charge. Our findings reveal the existence of tensor, vector, and scalar modes of gravitational waves. Firstly, the theory supports the + and $\times$ tensor modes propagating at the speed of light. Secondly, in the case where particles do not carry hypermomentum, vector modes propagating at the speed of light exist only within a very specific parameter space. However, when particles do carry hypermomentum, there are two shear modes that propagate at the speed of light, while the vector-$x$ and vector-$y$ modes emerge only under very specific conditions. Thirdly, in the presence of hypermomentum, there is always a longitudinal mode propagating at the speed of light. The universal existence of the shear modes and the longitudinal mode in the presence of hypermomentum is a key feature of symmetric teleparallel gravity, distinguishing it from the Riemannian framework through gravitational wave polarization detection. We also analyze the polarization modes in two widely studied special theories: $f(Q)$ theory and quadratic non-metricity theory. Our study reveals that, within the $f(Q)$ gravity framework, it is crucial to assume that matter fields are independent of the connection, as any dependence would lead to unphysical results.
@@ I Introduction
The recently proposed geometrical trinity of general relativity Lavinia Heisenberg1 ; Lavinia Heisenberg2 illustrates that, at the level of classical field equations, gravity theories equivalent to general relativity can be formulated within two geometric frameworks that differ fundamentally from Riemannian geometry. One such framework is metric teleparallel geometry Sebastian Bahamonde , which departs from Riemannian geometry by requiring that the spacetime curvature $R^{\mu}_{~{}\nu\lambda\rho}$ and the non-metricity $Q_{\lambda\mu\nu}$ vanish, while gravity arises purely from the torsion $T^{\lambda}_{~{}\mu\nu}$. The other framework, known as symmetric teleparallel geometry Lavinia Heisenberg3 , attributes gravity entirely to the non-metricity $Q_{\lambda\mu\nu}$, while both the curvature $R^{\mu}_{~{}\nu\lambda\rho}$ and the torsion $T^{\lambda}_{~{}\mu\nu}$ are set to zero.
@@
Since the field equations of these theories remain equivalent, there is no fundamental reason to favor Riemannian geometry over these alternative geometric descriptions of gravity. It is entirely conceivable that in an alternate parallel universe, an “Einstein” might have formulated a “general relativity” based on teleparallel gravity instead. Despite their formal equivalence, the distinct geometric foundations of these theories lead to different physical interpretations, offering deeper insights into the nature of gravity. Furthermore, their differences at the quantum level may provide new perspectives for tackling challenging issues in quantum gravity that remain unresolved within general relativity Lavinia Heisenberg4 ; Ruben Aldrovandi ; Lavinia Heisenberg5 . Finally, considering the modifications of gravity within these geometric frameworks can effectively bypass the constraints of Lovelock’s theorem David Lovelock1 ; David Lovelock2 , leading to new classes of modified gravity theories and providing novel approaches to addressing problems in quantum gravity and cosmology Sebastian Bahamonde ; Lavinia Heisenberg3 . For the above reasons, the geometrical trinity of general relativity and its modifications have received widespread attention and research.
@@
In this paper, we focus on various gravity theories within the symmetric teleparallel geometry framework and generally analyze the polarization modes of gravitational waves in these theories. The era of gravitational wave astronomy has arrived Abbott1 ; Abbott2 ; Abbott3 ; Abbott4 ; Abbott5 , which means that we can now and in the near future test various candidate modified gravity theories through high-precision gravitational wave detection J. Aasi ; F. Acernese ; T. Akutsu ; M. Punturo ; D. Reitze ; P. Amaro-Seoane ; Z. Luo ; Jun Luo ; Lijing Shao ; H. B. Jin ; Y. Y. Wang . Polarization, as a fundamental property of gravitational waves, is naturally one of the key focuses of gravitational wave detection. In four-dimensional spacetime, Ref. Eardley has already pointed out that under the assumption that matter fields couple only to the metric, a gravity theory can have at most six independent gravitational wave polarization modes. Furthermore, our previous research indicates that in a torsionless spacetime, test particles carrying hypermomentum charge may detect two additional polarization modes beyond the six known ones Yu-Qi Dong . Different theories predict different polarization modes of gravitational waves, so the detection of gravitational wave polarization can be used to test various gravity theories. This highlights the necessity of the research topic of this paper. There have already been extensive theoretical and technical studies on gravitational wave polarization modes Yu-Qi Dong ; f(R ; fT ; Orfeu Bertolami ; Horndeski0 ; TeVeS ; dCS and EdGB ; Y.Dong ; Y.Dong3 ; Y.Dong4 ; Y.Dong5 ; Y. Fan ; S. Bahamonde ; L.Shao ; Shaoqi Hou ; Xiao-Bin Lai ; Miguel Barroso Varela ; S Capozziello ; Kristen Schumacher1 ; Kristen Schumacher2 .
@@
Considering that in symmetric teleparallel gravity, the connection is not a priori expressed as a function of the metric, the existence of hypermomentum is naturally possible in the theory. Based on this, we consider two cases: one without hypermomentum and one with hypermomentum. In both cases, we analyze the polarization modes of gravitational waves in the most general symmetric teleparallel gravity that leads to second-order field equations. The analytical approach in this paper is based on the general framework developed in our previous research Y.Dong4 ; Y.Dong5 ; Yu-Qi Dong . This paper is organized as follows: In Sec. II, we discuss the linearized field equations of symmetric teleparallel gravity, which serve as the fundamental equations required for solving gravitational wave polarization modes. In Sec. III, we introduce the relevant concepts of gravitational wave polarization modes and explain their connection with gauge-invariant perturbations. These gauge-invariant perturbations significantly simplify the linearized field equations. In Sec. IV, we analyze the polarization modes of gravitational waves in general symmetric teleparallel gravity in two cases: when the test particle has hypermomentum and when it does not. In Sec. V, we specifically provide concrete examples of gravitational wave polarization modes in $f(Q)$ theory and quadratic non-metricity theory. These two theories have attracted significant attention and are representative. Finally, Sec. VI presents the conclusion.
@@ II Linearized symmetric teleparallel gravity
Here, $S_{g}$ is the action describing the gravitational part, while $S_{m}$ is the action describing the matter field $\Psi_{a}$. In the geometric framework of symmetric teleparallel gravity, there is no reason to require $S_{m}$ to be independent of the connection, thus naturally necessitating the consideration of the possible existence of hypermomentum. $\mathfrak{p}_{\mu}^{~{}\nu\lambda\rho}$ and $\mathfrak{q}_{\lambda}^{~{}\mu\nu}$ are Lagrange multipliers, which are tensors. Considering the symmetry of the indices in $R^{\mu}_{~{}\nu\lambda\rho}$ and $T^{\lambda}_{~{}\mu\nu}$, $\mathfrak{p}_{\mu}^{~{}\nu\lambda\rho}$ is antisymmetric with respect to the indices $(\lambda,\rho)$, while $\mathfrak{q}_{\lambda}^{~{}\mu\nu}$ is antisymmetric with respect to the indices $(\mu,\nu)$. As long as we note that the equation obtained by varying action (2) with respect to $\mathfrak{q}_{\lambda}^{~{}\mu\nu}$ only requires the torsion $T^{\lambda}_{~{}\mu\nu}=0$, and that the antisymmetric part of the indices $(\mu,\nu)$ in the equation obtained by varying action (2) with respect to the connection $\mathring{\Gamma}^{\lambda}_{~{}\mu\nu}$ only determines the value of the Lagrange multiplier $\mathfrak{q}_{\lambda}^{~{}\mu\nu}$, it can be proven that action (2) is equivalent to the following action:
@@
A good physical theory allows us to reasonably neglect the gravitational backreaction of test particles used for detecting gravitational waves on the spacetime background. Therefore, when studying the polarizations of gravitational waves, the vacuum case $S_{m}=0$ can always be considered, and the spacetime background can be set to flat. We assume that the flat spacetime background is
@@
The background equation obtained from the variation of the action with respect to the connection can be shown to satisfy
@@
To obtain the linear perturbation equations, it is sufficient to consider the second-order term in the perturbative expansion of action (3). By varying the second-order perturbation action with respect to the perturbations, we obtain the linear perturbation equations:
@@
This is because, by using Eqs. (5), (8), (9) and the symmetry of the indices, we can derive Eq. (11). In addition, for every solution $h_{\mu\nu}$ and $\Sigma^{\lambda}_{~{}\mu\nu}$ that satisfies Eqs. (10)-(12), there always exists $p_{\lambda}^{~{}\mu\nu\rho}$ (generally not unique, but this is not a problem) such that $h_{\mu\nu},\Sigma^{\lambda}_{~{}\mu\nu}$, and $p_{\lambda}^{~{}\mu\nu\rho}$ satisfy Eqs. (7)-(9). Equations (10)-(12) are the equations we need.
@@
Here, a series of $A,B,C$ with subscripts, such as $A_{(1)}$, is a set of constants. Action (13) can give us the most general forms of $\mathcal{F}_{\mu\nu}$ and $\mathcal{G}_{\lambda}^{~{}\mu\nu}$.
@@
We know that perturbations can be uniquely decomposed as Yu-Qi Dong ; Y.Dong6 ; James M. Bardeen ; Eanna E Flanagan ; Weinberg ; Evans
@@
where
@@
$h_{\mu\nu}\!\rightarrow\!h_{\mu\nu}-\partial_{\mu}\xi_{\nu}-\partial_{\nu}\xi_{\mu},\quad\Sigma^{\lambda}_{~{}\mu\nu}\!\rightarrow\!\Sigma^{\lambda}_{~{}\mu\nu}+\partial_{\mu}\partial_{\nu}\xi^{\lambda},$ (19)
@@
  • For transverse traceless third-order spatial tensor:
@@
  • For transverse traceless second-order spatial tensors:
@@
  • For transverse spatial vectors:
@@
That concludes all the necessary knowledge about gauge-invariant perturbations. Now, we turn to the fundamental concepts of gravitational wave polarization modes. The polarization modes of gravitational waves are measured by the relative motion of test particles. For the case without hypermomentum, the equation of relative motion between test particles is the well-known one:
@@
Figure 1: The six polarization modes of gravitational waves Eardley . Here, the gravitational wave propagates in the $+z$ direction. We have labeled the names of the six polarization modes in the figure. Based on the relationship between different polarization modes and gauge-invariant perturbations, the $+$ and $\times$ modes are referred to as tensor modes, the vector-$x$ and vector-$y$ modes are referred to as vector modes, and the breathing mode and longitudinal mode are referred to as scalar modes.
For the case with hypermomentum, the equation of relative motion for the test particle is Yu-Qi Dong
@@
The relationship between the polarization modes and gauge-invariant perturbations at this point is given by
@@ IV Gravitational wave polarization modes in the most general second-order symmetric teleparallel gravity
In this section, we analyze the gravitational wave polarization modes in general symmetric teleparallel gravity, which can yield second-order field equations. The equations we need to solve are Eqs. (10)-(12), where the explicit expressions for $\mathcal{F}_{\mu\nu}$ and $\mathcal{G}_{\lambda}^{~{}\mu\nu}$ are obtained from the variation of action (13).
@@ IV.1 Tensor modes
The form of the tensor equation is very simple. There is only one:
@@
Combining Eqs. (31) and (III), this implies that symmetric teleparallel gravity necessarily requires $A_{(2)}\neq 0$, and there must exist $+$ and $\times$ modes that propagate at the speed of light, regardless of whether the test particles carry hypermomentum charge.
@@ IV.2 Vector modes
There are three vector equations, and they are
@@
Since the equations are linear, they can be transformed into an algebraic system of equations by writing them in momentum space form for solving Yu-Qi Dong ; Y.Dong5 . The results show that, when $04A_{(2)}+C_{(1)}+E_{(1)}=$ 0, $\Xi_{i}$ propagates at the speed of light. According to Eqs. (31) and (III), this implies the existence of vector-$x$ and vector-$y$ modes that propagate at the speed of light. Under any other parameters, $\Xi_{i}=0$ and vector-$x$ and vector-$y$ modes do not exist. In particular, when the test particles do not carry hypermomentum charge, according to Eq. (31), vector mode gravitational waves appear only under a specific parameter choice.
As for the perturbation $K_{i}$, we find that when $C_{(1)}+E_{(1)}\neq 0$, there exists a solution $K_{i}$ that propagates at the speed of light. However, when $C_{(1)}+E_{(1)}=0$, $K_{i}$ is not constrained by any equation. For the case where test particles do not carry hypermomentum, the condition $C_{(1)}+E_{(1)}=0$ does not pose any issues within the range we have examined, because, according to Eq. (31), the unconstrained $K_{i}$ is not observable. However, according to Eq. (III), we find that $K_{i}$ contributes to shear modes. Therefore, if a symmetric teleparallel gravity allows the existence of hypermomentum, it must satisfy $C_{(1)}+E_{(1)}\neq 0$, otherwise, observable pathological gravitational wave polarization modes will exist. In addition, for the case of test particles carrying hypermomentum, shear modes that propagate at the speed of light always exist.

### v2

@@ Abstract
Abstract: In this paper, we investigate the polarization modes of gravitational waves within the most general symmetric teleparallel gravity theory that allows for second-order field equations We consider both scenarios where test particles either carry or do not carry a hypermomentum charge. Our findings reveal the existence of tensor, vector, and scalar modes of gravitational waves. Firstly, the theory supports the + and $\times$ tensor modes propagating at the speed of light. Secondly, in the case where particles do not carry hypermomentum, vector modes propagating at the speed of light exist only within a very specific parameter space. However, when particles do carry hypermomentum, there are two shear modes that propagate at the speed of light, while the vector-$x$ and vector-$y$ modes emerge only under very specific conditions. Thirdly, in the presence of hypermomentum, there is always a longitudinal mode propagating at the speed of light. The universal existence of the shear modes and the longitudinal mode in the presence of hypermomentum is a key feature of symmetric teleparallel gravity, distinguishing it from the Riemannian framework through gravitational wave polarization detection. We also analyze the polarization modes in two widely studied special theories: $f(Q)$ theory and quadratic non-metricity theory. Our study reveals that, within the $f(Q)$ gravity framework, it is crucial to assume that matter fields are independent of the connection, as any dependence would lead to unphysical results.
@@ I Introduction
The recently proposed geometrical trinity of general relativity Lavinia Heisenberg1 ; Lavinia Heisenberg2 ; Salvatore Capozziello Vittorio De Falco Carmen Ferrara ; Salvatore Capozziello Vittorio De Falco Carmen Ferrara2 illustrates that, at the level of classical field equations, gravitational theories equivalent to general relativity can be formulated within two geometric frameworks that differ fundamentally from Riemannian geometry. One such framework is metric teleparallel geometry Sebastian Bahamonde , which departs from Riemannian geometry by requiring that the spacetime curvature $R^{\mu}_{~{}\nu\lambda\rho}$ and the non-metricity $Q_{\lambda\mu\nu}$ vanish, while gravity arises purely from the torsion $T^{\lambda}_{~{}\mu\nu}$. The other framework, known as symmetric teleparallel geometry Lavinia Heisenberg3 , attributes gravity entirely to the non-metricity $Q_{\lambda\mu\nu}$, while both the curvature $R^{\mu}_{~{}\nu\lambda\rho}$ and the torsion $T^{\lambda}_{~{}\mu\nu}$ are set to zero.
@@
Since the field equations of these theories remain equivalent, there is no fundamental reason to favor Riemannian geometry over these alternative geometric descriptions of gravity. It is entirely conceivable that in an alternate parallel universe, an “Einstein” might have formulated a “general relativity” based on teleparallel gravity instead. Despite their formal equivalence, the distinct geometric foundations of these theories lead to different physical interpretations, offering deeper insights into the nature of gravity. Furthermore, their differences at the quantum level may provide new perspectives for tackling challenging issues in quantum gravity that remain unresolved within general relativity Lavinia Heisenberg4 ; Ruben Aldrovandi ; Lavinia Heisenberg5 . Finally, considering the modifications of gravity within these geometric frameworks can effectively bypass the constraints of Lovelock’s theorem David Lovelock1 ; David Lovelock2 , leading to new classes of modified gravitational theories and providing novel approaches to addressing problems in quantum gravity and cosmology Sebastian Bahamonde ; Lavinia Heisenberg3 . For the above reasons, the geometrical trinity of general relativity and its modifications have received widespread attention and research.
@@
In this paper, we focus on various gravitational theories within the symmetric teleparallel geometry framework and generally analyze the polarization modes of gravitational waves. The era of gravitational wave astronomy has arrived Abbott1 ; Abbott2 ; Abbott3 ; Abbott4 ; Abbott5 , which means that we can now and in the near future test various candidate modified gravitational theories through high-precision gravitational wave detection J. Aasi ; F. Acernese ; T. Akutsu ; M. Punturo ; D. Reitze ; P. Amaro-Seoane ; Z. Luo ; Jun Luo ; Lijing Shao ; H. B. Jin ; Y. Y. Wang . Polarization, as a fundamental property of gravitational waves, is naturally one of the key focuses of gravitational wave detection. In has been proven in Ref. Eardley that there are at most six independent gravitational wave polarization modes in a four-dimensional gravitational theory with matter fields coupling only to the metric. Furthermore, our previous research indicates that in a torsionless spacetime, test particles carrying hypermomentum charge may detect two additional polarization modes beyond the six known ones Yu-Qi Dong . Different theories predict different polarization modes of gravitational waves, so the detection of gravitational wave polarizations can be used to test various gravitational theories. This highlights the necessity of the research topic of this paper. There have already been extensive theoretical and technical studies of gravitational wave polarization modes Yu-Qi Dong ; f(R ; fT ; Orfeu Bertolami ; Horndeski0 ; TeVeS ; dCS and EdGB ; Y.Dong ; Y.Dong3 ; Y.Dong4 ; Y.Dong5 ; Y. Fan ; S. Bahamonde ; L.Shao ; Shaoqi Hou ; Xiao-Bin Lai ; Miguel Barroso Varela ; S Capozziello ; Kristen Schumacher1 ; Kristen Schumacher2 ; Ismail Soudi ; Manuel Hohmann ; Ligong Bian .
@@
Considering that in symmetric teleparallel gravity, the connection is not a priori expressed as a function of the metric, the existence of hypermomentum is naturally possible in the theory. Based on this, we consider two cases: one without hypermomentum and one with hypermomentum. In both cases, we analyze the polarization modes of gravitational waves in the most general symmetric teleparallel gravity that leads to second-order field equations. The analytical approach in this paper is based on the general framework developed in our previous research Y.Dong4 ; Y.Dong5 ; Yu-Qi Dong .
This paper is organized as follows: In Sec. II, we discuss the linearized field equations of symmetric teleparallel gravity, which serve as the fundamental equations required for solving gravitational wave polarization modes. In Sec. III, we introduce the relevant concepts of gravitational wave polarization modes and explain their connection with gauge-invariant perturbations. These gauge-invariant perturbations significantly simplify the linearized field equations. In Sec. IV, we analyze the polarization modes of gravitational waves in general symmetric teleparallel gravity. In Sec. V, we specifically provide concrete examples of gravitational wave polarization modes in $f(Q)$ theory and quadratic non-metricity theory. These two theories have attracted significant attention and are representative. Finally, Sec. VI presents the conclusion.
@@ II Linearized symmetric teleparallel gravity
Here, $S_{g}$ is the action describing the gravitational part, while $S_{m}$ is the action describing the matter field $\Psi_{a}$. In the geometric framework of symmetric teleparallel gravity, there is no reason to require $S_{m}$ to be independent of the connection, thus naturally necessitating the consideration of the possible existence of hypermomentum. $\mathfrak{p}_{\mu}^{~{}\nu\lambda\rho}$ and $\mathfrak{q}_{\lambda}^{~{}\mu\nu}$ are Lagrange multipliers, which are tensors. Considering the symmetry of the indices in $R^{\mu}_{~{}\nu\lambda\rho}$ and $T^{\lambda}_{~{}\mu\nu}$, $\mathfrak{p}_{\mu}^{~{}\nu\lambda\rho}$ is antisymmetric with respect to the indices $(\lambda,\rho)$, while $\mathfrak{q}_{\lambda}^{~{}\mu\nu}$ is antisymmetric with respect to the indices $(\mu,\nu)$. As long as we note that the equation obtained by varying the action (2) with respect to $\mathfrak{q}_{\lambda}^{~{}\mu\nu}$ only requires the torsion $T^{\lambda}_{~{}\mu\nu}=0$, and that the antisymmetric part of the indices $(\mu,\nu)$ in the equation obtained by varying the action (2) with respect to the connection $\mathring{\Gamma}^{\lambda}_{~{}\mu\nu}$ only determines the value of the Lagrange multiplier $\mathfrak{q}_{\lambda}^{~{}\mu\nu}$, it can be proven that the action (2) is equivalent to the following action:
@@
A good physical theory allows us to reasonably neglect the gravitational backreaction of test particles used for detecting gravitational waves on the spacetime background. Therefore, when studying the polarizations of gravitational waves, the vacuum case without matter can always be considered, and the spacetime background can be set to flat. We assume that the flat spacetime background is
@@
The background equations obtained from the variation of the action with respect to the connection can be shown to satisfy
@@
To obtain the linear perturbation equations, it is sufficient to consider the second-order terms in the perturbative expansion of the action (3). By varying the second-order perturbation action with respect to the perturbations, we obtain the linear perturbation equations:
@@
This is because, by using Eqs. (5), (8), (9) and the symmetry of the indices, we can derive Eq. (11). In addition, for every solution of $h_{\mu\nu}$ and $\Sigma^{\lambda}_{~{}\mu\nu}$ that satisfies Eqs. (10)-(12), there always exists $p_{\lambda}^{~{}\mu\nu\rho}$ (generally not unique, but this is not a problem) such that $h_{\mu\nu},\Sigma^{\lambda}_{~{}\mu\nu}$, and $p_{\lambda}^{~{}\mu\nu\rho}$ satisfy Eqs. (7)-(9). Equations (10)-(12) are the equations we need.
@@
Here, a series of $A,B,C$ with subscripts, such as $A_{(1)}$, is a set of parameters. The action (13) can give us the most general forms of $\mathcal{F}_{\mu\nu}$ and $\mathcal{G}_{\lambda}^{~{}\mu\nu}$.
@@
It is known that perturbations can be uniquely decomposed as Yu-Qi Dong ; Y.Dong6 ; James M. Bardeen ; Eanna E Flanagan ; Weinberg ; Evans
@@
where some constraints have been imposed:
@@
$h_{\mu\nu}\!\rightarrow\!h_{\mu\nu}-\partial_{\mu}\xi_{\nu}-\partial_{\nu}\xi_{\mu},\quad\Sigma^{\lambda}_{~{}\mu\nu}\!\rightarrow\!\Sigma^{\lambda}_{~{}\mu\nu}+\partial_{\mu}\partial_{\nu}\xi^{\lambda}.$ (19)
@@
  • Transverse traceless third-order spatial tensor:
@@
  • Transverse traceless second-order spatial tensors:
@@
  • Transverse spatial vectors:
@@
This concludes all the necessary knowledge about gauge-invariant perturbations. Now, we turn to the fundamental concepts of gravitational wave polarization modes. The polarization modes of gravitational waves are measured by the relative motion of test particles. For the case without hypermomentum, the equation of relative motion between two test particles is the well-known one:
@@
Figure 1: The six polarization modes of gravitational waves Eardley . Here, the gravitational wave propagates in the $+z$ direction. We have labeled the names of the six polarization modes in the figure. Based on the relationship between different polarization modes and gauge-invariant perturbations, the $+$ and $\times$ modes are referred to as tensor modes, the vector-$x$ and vector-$y$ modes are referred to as vector modes, and the breathing and longitudinal modes are referred to as scalar modes.
For the case with hypermomentum, the equation of relative motion for the test particles is Yu-Qi Dong
@@
The relationship between the polarization modes and gauge-invariant perturbations in this case is given by
@@ IV Gravitational wave polarization modes in the most general second-order symmetric teleparallel gravity
In this section, we analyze the gravitational wave polarization modes in general symmetric teleparallel gravity, which can yields second-order field equations. The equations we need to solve are Eqs. (10)-(12), where the explicit expressions for $\mathcal{F}_{\mu\nu}$ and $\mathcal{G}_{\lambda}^{~{}\mu\nu}$ are obtained from the variation of the action (13).
@@ IV.1 Tensor modes
The equation for the tensor perturbation is given by
@@
Combining Eqs. (31) and (III), this implies that symmetric teleparallel gravity necessarily requires $A_{(2)}\neq 0$, and there must exist $+$ and $\times$ modes propagating at the speed of light, regardless of whether the test particles carry hypermomentum charge.
@@ IV.2 Vector modes
There are three vector equations:
@@
Since these equations are linear, they can be transformed into an algebraic system by rewriting them in momentum space form to facilitate solving Yu-Qi Dong ; Y.Dong5 . The results show that, when $04A_{(2)}+C_{(1)}+E_{(1)}=$ 0, $\Xi_{i}$ propagates at the speed of light. According to Eqs. (31) and (III), this implies the existence of vector-$x$ and vector-$y$ modes that propagate at the speed of light. For any other parameters, $\Xi_{i}=0$ and vector-$x$ and vector-$y$ modes do not exist. In particular, when the test particles do not carry hypermomentum charge, vector mode gravitational waves emerge solely within a specific parameter space, as dictated by Eq. (31).
We find that when $C_{(1)}+E_{(1)}\neq 0$, there exists a solution for the perturbation $K_{i}$ and it propagates at the speed of light. However, when $C_{(1)}+E_{(1)}=0$, $K_{i}$ is not constrained by any equation. For the case where test particles do not carry hypermomentum, the condition $C_{(1)}+E_{(1)}=0$ does not pose any issues within the range we have examined, because, according to Eq. (31), the unconstrained $K_{i}$ is not observable. However, according to Eq. (III), we find that $K_{i}$ contributes to shear modes. Therefore, if a symmetric teleparallel gravity allows the existence of hypermomentum, it must satisfy $C_{(1)}+E_{(1)}\neq 0$, otherwise, observable pathological gravitational wave polarization modes will exist. In addition, for the case of test particles carrying hypermomentum, shear modes that propagating at the speed of light always exist.